# MineInsight: A Multi-sensor Dataset for Humanitarian Demining Robotics in Off-Road Environments

### v1

@@ Abstract
The use of robotics in humanitarian demining increasingly involves computer vision techniques to improve landmine detection capabilities. However, in the absence of diverse and realistic datasets, the reliable validation of algorithms remains a challenge for the research community. In this paper, we introduce MineInsight, a publicly available multi-sensor, multi-spectral dataset designed for off-road landmine detection. The dataset features 35 different targets (15 landmines and 20 commonly found objects) distributed along three distinct tracks, providing a diverse and realistic testing environment. MineInsight is, to the best of our knowledge, the first dataset to integrate dual-view sensor scans from both an Unmanned Ground Vehicle and its robotic arm, offering multiple viewpoints to mitigate occlusions and improve spatial awareness. It features two LiDARs, as well as images captured at diverse spectral ranges, including visible (RGB, monochrome), visible short-wave infrared (VIS-SWIR), and long-wave infrared (LWIR). Additionally, the dataset comes with an estimation of the location of the targets, offering a benchmark for evaluating detection algorithms. We recorded approximately one hour of data in both daylight and nighttime conditions, resulting in around 38,000 RGB frames, 53,000 VIS-SWIR frames, and 108,000 LWIR frames. MineInsight serves as a benchmark for developing and evaluating landmine detection algorithms. Our dataset is available at https://github.com/mariomlz99/MineInsight.
@@ I INTRODUCTION
We provide an estimation of target locations, which include 3D positions on a global map and corresponding 2D bounding boxes in camera frames, serving as a benchmark for validating detection algorithms and facilitating further research and development in the field.
@@ II RELATED WORK
Platform | Environment | Camera | LiDAR | Labels | ||||||||||||
@@
Binghamton University has contributed significantly to the field of landmine detection by releasing multiple datasets as part of the Scatterable Landmine Detection Project. These datasets include data from various UAV flights flown over different terrains. The project’s primary focus is the detection of PFM-1 landmines, together with their carrying case. We propose a categorization of the dataset releases into three groups (1–7, 8–9, and 10–24), facilitating a direct comparison between UAV platforms, sensors, and data collection methods across studies. The first group (Datasets 1-7), released by De Smet et al. [12], includes data collected using a DJI Phantom 4 Pro UAV and a DJI Matrice 600 Pro UAV, both equipped with a Parrot Sequoia Multispectral Camera. Flights were conducted over grass and rubble environments, resulting in seven high-resolution orthophotos. Steinberg et al. released the second group (Datasets 8-9) [13], which includes data collected from three flights over grass and four flights over rubble using a DJI Matrice 600 Pro UAV equipped with a Parrot Sequoia Multispectral Camera. Baur et al. later used the same data to explore the potential of deep learning-based landmine detection, with their experimental setup well described in [3], where they address the challenges of detecting PFM-1 landmines in UAV-based imagery. Baur et al. [14] released the third group (Datasets 10-24), which comprises photogrammetric data collected using a DJI Phantom 4 UAV equipped with a Parrot Sequoia Multispectral Camera and processed with Pix4DMapper. To introduce variability, the PFM-1 landmines were rotated, partially buried, or partially hidden in each orthophoto.
@@ IV DATASET
In this section, we present the MineInsight dataset [25]. We describe the targets, detail the environmental conditions with varying vegetation and lighting, and outline the recorded data sequences, including the UGV and robotic arm configurations. Finally, we explain the automated reference data generation process.
@@ IV-A Targets
We placed a combination of AP and AT inert landmines and common items that might be found in natural environments across three tracks to construct a diverse dataset. The items were left overnight, allowing their temperatures to equilibrate with the ambient environment and better mimic realistic conditions. All the items included in the dataset are shown in Figure 2. The following is an overview of the variety of landmines and common items included per track.
Track 1 (25 targets):
Landmines: 9x PFM-1 “Butterfly”(AP), 2x PMN (AP), M6 (AT), TMA-2 (AT), TC-3.6 (AT);
Common items: A plastic cup, a metal tuna can, a metal coke can, a plastic water bottle (significantly crumpled), a plastic water bottle (partially crumpled), a metal pot, a glass beer bottle, a plastic chips bag (significantly crumpled), a glass pepper dispenser, and a disposable paper cup.
Track 2 (21 targets):
Landmines: 6x PFM-1 (AP), MON-90 (AT), Type 72P (AT), TM-46 (AT), TMM-1 (AT), MON-50 (AT);
Common items: A plastic water bottle (slightly crumpled), a sponge, a plastic charger, a plastic chips bag (slightly crumpled), a glass jar (green cover), a soda metal can, a glass jar (grey cover), a glass vinegar bottle, a metal corn tin, and a plastic cup.
Track 3 (21 targets):
Landmines: 6x PFM-1 (AP), C-3 “Elsie” (AP), PROM-1 (AP), M-36 (AP), VS-50 (AT);
Common items: A glass jar (green cover), a glass jar (grey cover), a sponge, a plastic charger, a plastic cup, a plastic chips bag (slightly crumpled), a glass vinegar bottle, a soda metal can, a metal corn tin, and a plastic water bottle (slightly crumpled).
@@ IV-B Environment
We selected three outdoor environments for data sampling, each consisting of an approximately long and wide track designated for inspection. The track boundaries are marked by ropes to indicate restricted access. The terrain is primarily covered with leaves. Temperatures during the test campaign ranged from to under consistently cloudy conditions (no direct sunlight), but these temperature measurements are not included in the dataset. We categorize the tracks as follows:
Track 1: Low to mid-high vegetation, data collected between 1:00 PM and 2:00 PM. Visible light, temperature around ;
Track 2: Low to mid-high vegetation, data collected between 3:00 PM and 4:00 PM. Visible light, temperature around ;
Track 3: Mid-to-high vegetation, with denser and taller growth compared to the previous tracks, data collected between 5:00 PM and 6:00 PM. Recorded in dark conditions, temperature around . The RGB camera is unavailable for this track primarily due to a hardware failure, although the low light would also have limited its usefulness.
An overview of the vegetation conditions for each track, as well as the distribution of landmines and common items, is provided in Figure 3.
@@ IV-C Sequences
For each lane, we recorded three sequences: the reference sequence and two evaluation sequences. However, only the evaluation sequences are included in the released dataset. We use the reference sequence only for reference data generation, as detailed in section IV-D. The first and second evaluation sequences, referred to as sequence 1 and sequence 2, differ in the position of the robotic arm during data sampling.
@@ IV-D Reference data generation
As illustrated in Figure 6, targets can hardly be detected by the human eye in the images. Conventional data labeling methods are impractical for our images due to the low visibility of targets. Therefore, we implement an automated labeling system based on AprilTags [26], a LiDAR-inertial SLAM system [27], and the relative pose of the sensors with respect to the robot base (incorporating the relevant kinematic transformations). We placed an AprilTag at each target location on every track and recorded a reference sequence. An example of a track prepared for the reference sequence, along with a sample target and its corresponding tag, is shown in Figure 5. The main steps of the reference data generation process are the following:
@@
Finally, using the intrinsic and extrinsic calibration parameters of the respective camera, we project the transformed target locations onto the image plane to obtain the pixel coordinates of the center of each target.
V CONCLUSION
We introduced MineInsight, a multi-sensor, multi-spectral dataset designed to advance robotic landmine detection in off-road environments. Our dataset combines dual-sensor views (mobile base and robotic arm), multi-spectral imaging (monochrome, RGB, VIS-SWIR, LWIR), dual LiDAR sensors, and synchronized UGV and arm data to overcome occlusions and sensor limitations in demining operations. We believe that this dataset is well-suited for validating landmine detection algorithms. Moreover, given the rich diversity of the off-road data, we encourage researchers to use the dataset for generic off-road robotics tasks, such as navigation and mapping, alongside its primary application in demining operations. Our future work will exploit spectral fusion or landmine segmentation to further enhance sensing for demining. In addition, we plan to expand the dataset by adding new tasks, upgrading hardware (including additional sensors), incorporating diverse scenarios (such as different seasons or varying weather conditions), and including additional landmine types.
@@ ACKNOWLEDGMENT
The authors thank Alessandra Miuccio and Timothée Fréville for their support in the hardware and software design. They also thank Sanne Van Hees and Jorick Van Kwikenborne for their assistance in organizing the measurement campaign.

### v2

@@ Abstract
The use of robotics in humanitarian demining increasingly involves computer vision techniques to improve landmine detection capabilities. However, in the absence of diverse and realistic datasets, the reliable validation of algorithms remains a challenge for the research community. In this paper, we introduce MineInsight, a publicly available multi-sensor, multi-spectral dataset designed for off-road landmine detection. The dataset features 35 different targets (15 landmines and 20 commonly found objects) distributed along three distinct tracks, providing a diverse and realistic testing environment. MineInsight is, to the best of our knowledge, the first dataset to integrate dual-view sensor scans from both an Unmanned Ground Vehicle and its robotic arm, offering multiple viewpoints to mitigate occlusions and improve spatial awareness. It features two LiDARs, as well as images captured at diverse spectral ranges, including visible (RGB, monochrome), visible short-wave infrared (VIS-SWIR), and long-wave infrared (LWIR). Additionally, the dataset provides bounding boxes generated by an automated pipeline and refined with human supervision. We recorded approximately one hour of data in both daylight and nighttime conditions, resulting in around 38,000 RGB frames, 53,000 VIS-SWIR frames, and 108,000 LWIR frames. MineInsight serves as a benchmark for developing and evaluating landmine detection algorithms. Our dataset is available at https://github.com/mariomlz99/MineInsight.
@@ I Introduction
We provide 2D bounding boxes refined with human supervision, serving as a benchmark for validating detection algorithms and supporting further research.
@@ II Related Work
Platform | Environment | Camera | LiDAR | Labels | |||||||||||
@@
Binghamton University has contributed significantly to the field of landmine detection by releasing multiple datasets as part of the Scatterable Landmine Detection Project. These datasets include data from various UAV flights flown over different terrains. The project’s primary focus is the detection of PFM-1 landmines, together with their carrying case. We propose a categorization of the dataset releases into three groups (1–7, 8–9, and 10–24), facilitating a direct comparison between UAV platforms, sensors, and data collection methods across studies. The first group (Datasets 1-7), released by De Smet et al. [6], includes data collected using a DJI Phantom 4 Pro UAV and a DJI Matrice 600 Pro UAV, both equipped with a Parrot Sequoia Multispectral Camera. Flights were conducted over grass and rubble environments, resulting in seven high-resolution orthophotos. Steinberg et al. released the second group (Datasets 8-9) [23], which includes data collected from three flights over grass and four flights over rubble using a DJI Matrice 600 Pro UAV equipped with a Parrot Sequoia Multispectral Camera. Baur et al. later used the same data to explore the potential of deep learning-based landmine detection, with their experimental setup described in [4], where they address the challenges of detecting PFM-1 landmines in UAV-based imagery. Baur et al. [2] released the third group (Datasets 10-24), which comprises photogrammetric data collected using a DJI Phantom 4 UAV equipped with a Parrot Sequoia Multispectral Camera and processed with Pix4DMapper. To introduce variability, the PFM-1 landmines were rotated, partially buried, or partially hidden in each orthophoto.
@@ IV Dataset
In this section, we present the MineInsight dataset [15]. We describe the targets, detail the environmental conditions with varying vegetation and lighting, and outline the recorded data sequences, including the UGV and robotic arm configurations. Finally, we explain the reference data generation process, which combines an automated pipeline with human supervision to improve the quality of the annotations.
@@ IV-A Targets
We placed a combination of AP and AT inert landmines and common items that might be found in natural environments across three tracks to construct a diverse dataset. The items were left overnight, allowing their temperatures to equilibrate with the ambient environment and better mimic realistic conditions. All the items included in the dataset are shown in Figure 2. We summarize the distribution of landmines and common items per track as follows:
Track 1 (25 targets): 15 landmines (11 AP, 4 AT) and 10 common items. Track 2 (21 targets): 11 landmines (6 AP, 5 AT) and 10 common items. Track 3 (21 targets): 11 AP landmines and 10 common items.
@@ IV-B Environment
We selected three outdoor environments for data sampling, each consisting of an approximately long and wide track designated for inspection. The track boundaries were marked by ropes to indicate restricted access. The terrain was primarily covered with leaves, and the campaign took place under consistently cloudy conditions (no direct sunlight). We categorize the tracks as follows:
Track 1: Forest floor with low to mid-high vegetation, mostly covered by freshly fallen deciduous leaves over moist clay-loam soil. Recorded between 13:00 and 14:15, with ambient air temperatures ranging from to . Conditions were overcast, giving diffuse light and a uniform thermal background;
Track 2: Low to mid-high vegetation, with a more heterogeneous leaf litter than Track 1. The surface shows uneven layers of leaves with occasional gaps exposing the moist clay-loam soil beneath. Recorded between 15:15 and 16:05, with ambient air temperatures between and . Moisture variations created visible contrasts across modalities, with thermal differences between leaf-covered and exposed ground;
Track 3: Denser and taller vegetation compared to the previous tracks, with soil partly visible under uneven leaf cover. Recorded between 17:30 and 18:00, under twilight conditions, with ambient air temperatures between and . Cooling soil and foliage created mixed thermal patterns. The RGB camera is unavailable for this track due to a hardware failure, although the twilight conditions would have limited its usefulness.
An overview of the vegetation conditions for each track, as well as the distribution of landmines and common items, is provided in Figure 3. In addition, we release a minute-by-minute climatology log covering both the target placement phase and the campaign itself. The log records atmospheric variables (temperature, humidity, wind, pressure, cloud coverage, precipitation) as well as soil temperatures at multiple depths, providing the environmental context needed to interpret the thermal data.
@@ IV-C Sequences
For each lane, we recorded three sequences: the reference sequence and two evaluation sequences. We use the reference sequence only for reference data generation, as detailed in section IV-D. The first and second evaluation sequences, referred to as Sequence 1 and Sequence 2, differ in the position of the robotic arm during data sampling.
@@ IV-D Reference data generation
As illustrated in Figure 6, targets can hardly be detected by the human eye in the images. Direct manual labeling is difficult, as many targets are barely visible in the raw imagery. Therefore, we implement an annotation pipeline that generates initial estimates using AprilTags [19], SLAM [18], ICP, and sensor kinematics, which are then adjusted by human intervention. We placed an AprilTag at each target location on every track and recorded a reference sequence. An example of a track prepared for the reference sequence, along with a sample target and its corresponding tag, is shown in Figure 5. The main steps of the reference data generation process are the following:
@@
Finally, using the intrinsic and extrinsic calibration parameters of the respective camera, we project the transformed target locations onto the image plane to obtain the pixel coordinates of the center of each target, and apply a fixed-size bounding box according to the corresponding camera resolution.
5.
Human-supervised adjustment of bounding boxes: The automatically generated boxes serve as an initial guess. Annotators adjust them to better match the visible extent of each target, mitigating errors propagated from AprilTag detection, pose estimation, or perspective distortions. RGB and VIS–SWIR annotations are manually revised. For LWIR, where signatures are often faint, adjustments rely on reprojections from the corrected RGB and VIS–SWIR annotations, with human input.
V Baseline Evaluation
As a baseline, we trained YOLOv8 [10] on the SULAND dataset [27], which is the closest publicly available dataset to our scenario. Even with extensive augmentation strategies, the model did not transfer successfully to MineInsight and failed to produce reliable detections. Rather than serving as a performance benchmark, this outcome highlights the domain gap between existing datasets and MineInsight, driven by its clutter-rich environments with vegetation and distractor objects, as well as a broader variety of landmine types. This underlines MineInsight’s value as a benchmark for studying domain adaptation and multi-modal fusion in realistic demining scenarios.
VI Limitations
MineInsight is, to the best of our knowledge, the first multi-modal dataset for landmine detection with RGB, VIS–SWIR, and LWIR sensing. Its current release, however, has some constraints. All data were collected during a single late-autumn campaign under overcast conditions. As a result, models trained on this dataset may face domain gaps when applied to different seasons, terrains, or weather conditions, such as sunny, snowy, arid, or grassy environments. The release also lacks RGB data for Track 3, though twilight conditions would in any case have limited its usefulness. While MineInsight surpasses existing public datasets in scale and sensor diversity, the number of targets and total recording time remain modest relative to real-world scenarios. More broadly, the dataset highlights the domain-adaptation challenges faced when transferring models trained elsewhere to MineInsight. Finally, annotations are adjusted with human supervision; minor inaccuracies may remain, particularly for faint thermal signatures. These limitations motivate the future extensions outlined in the conclusion.
VII Conclusion
We introduced MineInsight, a multi-sensor, multi-spectral dataset for robotic landmine detection in off-road environments. It combines dual viewpoints (UGV and robotic arm), multi-spectral imaging (monochrome, RGB, VIS–SWIR, LWIR), dual LiDARs, and synchronized data to address occlusions and sensor limitations. The dataset includes bounding boxes generated by an automated pipeline with a human-in-the-loop process to support the evaluation of detection algorithms, and it can also serve generic off-road robotics tasks such as navigation and mapping. In addition, the dataset includes minute-by-minute climatology measurements, which provide valuable context for exploiting LWIR data and studying thermal contrast under varying conditions. As noted in Section VI, the current release is limited to a single late-autumn campaign and may introduce domain gaps under other environmental conditions. Future work will expand MineInsight with recordings across varied terrains, climates, and seasons, additional sensors, new tracks, and further landmine types. We also aim to explore spectral fusion, landmine segmentation, and synthetic data for domain adaptation. Finally, the issue of RGB degradation at night, observed in Track 3, could be further studied with lighting systems to extend visibility.
@@ ACKNOWLEDGMENT
The authors thank Alessandra Miuccio and Timothée Fréville for their support in the hardware and software design. They also thank Sanne Van Hees and Jorick Van Kwikenborne for their assistance in organizing the measurement campaign. Finally, they thank the Belgian Meteo Wing for providing the climatology study during the days of the test campaign.